The ratio of total surface area to lateral surface area of cylinder whose radius is 80cm and height 20cm is?

Grade:8

2 Answers

kavin bala s
6 years ago
TSA=(2πr*r+2πrh).LSA= 2πrh.So,TSA/LSA=(2πr*r+2πrh)2πh=1+(80/20)=5. If it not clears your doubt pls reply to me.
Sahil Balyan
6 years ago
r=80cm, h=20cmTSA=2πr(h+r)TSA=2π(80)(20+80)TSA=2π(80)(100)............ (1)LSA=2πrhLSA=2π(80)(20)..........(2)eq. (1)/eq.(2)We getTSA/LSA=16000π/3200π=5so, TSA:LSA=5:1